OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = −x2 + 2.5x + 1. What do the solutions of 0 = −x2 + 2.5x + 1 represent on the graph of the equation?

OpenStudy (anonymous):

Take a look at your equation: 0=−x2+2.5x+1 This should look suspiciously similar to this: y=−x2+2.5x+1 The reason for that is that your equation is a special case of the one I just showed you, for the case of y = 0. But y=0 is the x-axis! That means that if you graph the equation, the points where the graph crosses the x-axis are the solutions to the equation.

OpenStudy (anonymous):

that doesnt really answer the question doe

OpenStudy (anonymous):

how

OpenStudy (anonymous):

just read the question and then what you wrote

OpenStudy (anonymous):

here hold on

OpenStudy (anonymous):

The solutions would represent the x-intercepts on the graph, since the y-value is zero. Using a ti-84, you can simply graph y=-x^2+2.5x+1, and one of the options around the top 5 buttons (probably calc) will give you a menu that you can scroll down. One of the options will allow you to calculate the x-intercept. However, you should check if the solutions are integers first. Press 2nd, then graph, which will show you the table. Look for the x value when the corresponding y value is 0.

OpenStudy (anonymous):

hahaha whattt??

OpenStudy (anonymous):

its hard to understand

OpenStudy (anonymous):

yeahh, you dont say?

OpenStudy (anonymous):

i gtg i will be back on in fourth block ok babe by i love u

OpenStudy (anonymous):

byeee

OpenStudy (anonymous):

IM BACKK!!!

OpenStudy (anonymous):

I SEE!! HAHAH

OpenStudy (anonymous):

BABE???? is that your boyfriend?

OpenStudy (anonymous):

ya ii am her bf

OpenStudy (anonymous):

ok